I have an opportunity to invest $28,000 today at a compound interest rate of 7.1% p.a compounded monthly. I intend to withdraw the total account balance when it reaches $118,000. How long do I need to hold the investment? Answer in months to two decimal places.
Answer #1

Future value = present value (1+r)^n

Where, r is interest rate and n is monthly time.

Monthly interest rate = 7.1/12 = 0.5916%

1. 18,000 = 28000(1 +0.005916)

4.21429 = (1.005916)

(1.0059161/244) = (1.005916)

n = 244

It will take 244 months approximately to reach investment at $118000.